SD meane 3172.42 L 413.8 guzzlere Non-guzzler Sample size 582 9852 1917.32 484.52Intellectual Property of Katie Jager © Calculate and Interpret Output: (5 points) e. (2 points) From the summary statistics c
Answer #1

we have given two sample :- & consider sample 1 = xi guzzler sample 2 = 72 = Non -94221 er. for x, n, - sample size = 58 samp2,. 1914.3 - Sample mean = 484.5. S2 = sample standard deviahon *) calculate & Interpret the output :- e) from the summary St11 1255.0 1255:1 56.4851 3190.5639 = 22.2200 degree of freedom. nith 2-2 (Assume equal variance) = 58+985-2 1041 9 Р value obb 99% confidence interval for cl, ~42) for x, f where N,= Population mean M₂= Population mean for a 2 (H1-H2) =( cat_S2) I tatool + 2.5805 (21904 Whey) Tholi (M.-H2) =( 1255. + (2.5805 *56.4851)] (1255-1 I 145.7598] -(1109.340, 1400.8598) Interpretatoptain e value & confidence interval using R. e value= 2pt (-abs (t) df) (for two failed) f Pt (-abs (t), df) (for one tailed> ## p value from t test : 1 > 2*pt (-abs (22.22) [1] 7.80193e-90 1041) : (u1 u2) > ## Confidence interval > n1= 58 > x1bar =

## answer fro p value and confidence interval are same from by hand and r software , only there is difference in decimal points after answers